Resources Hint 12 of 25 > What is the mass, in grams, of 9.70 x 1024 molecules of methanol (CH,OH) ? mass: - a $ ote
Answer #1

1 mole of a substance contains an Avogadro number of molecules

So, Number of moles = number of molecules / Avogadro number = 9.70*1024 / 6.022*1023 = 16.1

The molar mass of methanol = atomic mass of C + 4* atomic mass of H + atomic mass of O

= 12+4*1+16 = 32 g / mole

Mass of methanol = number of moles * molar mass = 16.1*32 = 515.44 g